Circular No. 7199

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 L2 M. Elowitz and F. Shelly have reported the discovery of a comet with

a coma but no distinct tail in LINEAR data. In response to posting on The NEO Confirmation Page, G. R. Viscome (Lake Placid, NY) reports that the object shows a 16′′ coma and moderately strong condensation, but again no discernible tail. The following observations are available:

1999 UT α_{2000} δ_{2000} Mag. Observer
June 11.24296 16^{h}35^{m}06.^{s}96 +15^{o}25′47′′.4  18.6 LINEAR
11.25596 16 35 06.01 +15 25 23.6  18.9  _{”}
11.26886 16 35 05.05 +15 24 59.4  19.1  _{”}
11.29478 16 35 03.06 +15 24 14.3  19.0  _{”}
12.19718 16 33 58.91 +14 57 15.5  16.5 Viscome
12.20328 16 33 58.47 +14 57 04.6  _{”}
12.21365 16 33 57.71 +14 56 45.3  16.5  _{”}

SUPERNOVAE 1999ch AND 1999ci A. Gal-Yam and D. Maoz, Tel Aviv University; and R. Pogge, Ohio

State University, communicate: “Spectrograms (range 360–700 nm) of SNe 1999ch and 1999ci (cf. IAUC 7181), obtained by Pogge with the Mayall 4-m telescope (+ RC spectrograph) at Kitt Peak on June 8–9, confirm these to be type-Ia supernovae. SN 1999ch is ∼ 4 weeks past maximum brightness, at a redshift of z = 0.15, consistent with that of Abell 2235 (z = 0.151). SN 1999ci is ∼ 2 weeks past maximum, at a redshift of z = 0.12, consistent with that of Abell 1984 (z = 0.124).”

COMET C/1998 T1 (LINEAR) Visual m_{1} estimates: May 23.88 UT, 11.9 (A. Pearce, Nedlands, W.

Australia, 0.41-m reflector); June 8.67, 10.7 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector); 9.72, 10.3 (S. T. Rae, Palmerston North, New Zealand, 0.25-m reflector); 10.65, 10.2 (Rae).

COMET 29P/SCHWASSMANN-WACHMANN 1 Visual m_{1} estimates: Apr. 12.54 UT, 12.8 (D. A. J. Seargent, The En-

trance, N.S.W., 0.25-m reflector); May 18.33, 13.8 (R. Keen, Mt. Thorodin, CO, 0.32-cm reflector); 22.72, 13.9 (Pearce); June 12.21, 12.9 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1999 June 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green